### v3

# Tera-MIND: Tera-scale mouse brain simulation via spatial mRNA-guided diffusion

Tera-MIND: Tera-scale mouse brain simulation via spatial mRNA-guided diffusion
Jiqing Wu Department of Biomedical Engineering, University of Basel, Switzerland Ingrid Berg Department of Pathology and Molecular Pathology, University Hospital, University of Zurich, Switzerland Yawei Li Computer Vision Lab, ETH Zurich, Switzerland Integrated System Laboratory, ETH Zurich, Switzerland Ender Konukoglu Computer Vision Lab, ETH Zurich, Switzerland Viktor H. Koelzer Department of Biomedical Engineering, University of Basel, Switzerland Institute of Medical Genetics and Pathology, University Hospital Basel, Switzerland Correspondence: Viktor.Koelzer@usb.ch
SUMMARY
Holistic 3D modeling of molecularly defined brain structures is crucial for understanding complex brain functions. Using emerging tissue profiling technologies, researchers charted comprehensive atlases of mammalian brain with sub-cellular resolution and spatially resolved transcriptomic data. However, these tera-scale volumetric atlases pose computational challenges for modeling intricate brain structures within the native spatial context. We propose Tera-MIND, a novel generative framework capable of simulating Tera-scale Mouse braINs in 3D using a patch-based and boundary-aware Diffusion model. Taking spatial gene expression as conditional input, we generate virtual mouse brains with comprehensive cellular morphological detail at teravoxel scale. Through the lens of 3D gene-gene self-attention, we identify spatial molecular interactions for key transcriptomic pathways, including glutamatergic and dopaminergic neuronal systems. Lastly, we showcase the translational applicability of Tera-MIND on previously unseen human brain samples. Tera-MIND offers an efficient generative modeling of whole virtual organisms, paving the way for integrative applications in biomedical research.
KEYWORDS
Spatial transcriptomics, mouse and human brain, Generative AI, diffusion model
INTRODUCTION
Recently, a network of researchers 1, 2, 3, 4 has reached a groundbreaking milestone by charting the first complete cellular atlas of the adult mouse brain, the most extensively studied animal model in neuroscience. Using advanced spatial transcriptomics methods 5, 6, 7 capable of profiling molecularly defined brain anatomy and morphology at sub-cellular resolution, the collective studies established foundational resources and datasets, providing an unprecedented framework for systematically exploring the molecular-driven spatial organization of the mammalian brain. Consequently, this landmark achievement opens the door for next-generation investigations into intricate brain functions characterized by exceptional spatial complexity. However, these tera-scale volumetric datasets, including spatial molecular data (e.g., mRNA readouts acting as proxies for gene expression) and paired morphological bioimages, pose substantial computational challenges 8 for in-silico modeling of the mouse brain. Integrating spatial gene expression with 3D brain morphology to uncover functional relationships at whole organ scale remains an open challenge.
Meanwhile, generative artificial intelligence (GenAI) has emerged as a transformative tool for automating visual content creation and manipulation 9, 10. Given a text prompt that describes the desired visual content, GenAI models 11, 12 are able to produce highly realistic images and videos 13, 14 with remarkable efficiency. In biomedical research, the applications of GenAI have gained domain-specific recognition in various areas such as digital histopathology 15, drug screening 16, 17, and molecule design 9. However, the direct application of existing GenAI methods to whole organ simulation has currently remained infeasible due to technological limitations. Developed upon the framework of Generative Adversarial Nets (GAN) 11, the GigaGAN 18 model was trained on 96-128 A100 GPUs to synthesize $4,096\times 4,096$ color images. Using up to 6,144 H100 GPUs, Movie Gen 19, a diffusion-based model 12, achieved video generation results with $1,092\times 1,080\times 256$ spatial-temporal resolution. By contrast, each whole slide image (WSI; $0.108\mu m/\mathrm{pixel}$) of the quality-controlled mouse brain atlas has a spatial resolution of $73,216\times 105,984$. Consequently, stacking 50 slices of these DAPI (4’,6-diamidino-2-phenylindole) and PolyT-stained (total mRNA signals) WSIs results in $0.77\times 10^{12}$ voxels in total. This immense data scale of WSIs and paired gene expression arrays significantly intensifies the computational demands associated with GenAI modeling of a virtual mouse brain.
Addressing memory bottlenecks in high-resolution image generation has previously motivated the development of patch-based approaches 20, 21, 22, 23, 24, which enforce boundary consistency between generated neighboring patches. Similar to many standard GenAI models, these approaches can also leverage the interplay between textual and visual modalities and create desired visual content conditioned on text prompts. While this design choice has shown remarkable success in natural image synthesis and artistic content creation, it is suboptimal for biomedical applications – particularly for the accurate 3D molecular-driven reconstruction of mammalian brains. The inherent complexity of brain morphology controlled by spatial molecular interactions demands a generative framework that can capture molecular-to-morphology spatial associations with precision. In response to these challenges, we propose Tera-MIND, a novel GenAI approach designed for the paired data structure of spatial mRNA arrays (gene expression, prompts) and WSIs (brain morphology, bioimages). Specifically, Tera-MIND employs a newly introduced patch-based and boundary-aware diffusion model, which allows the reconstruction of tera-scale mouse brains with high fidelity. Owing to the patch-wise training and inference paradigm, Tera-MIND models in-silico mammalian brain(s) with computational efficiency. As a result, the whole simulation process can be efficiently executed on a single DGX A100 machine, significantly lowering hardware demands while maintaining scalability for tera-scale data processing. This efficiency underscores the potential wide applicability of Tera-MIND for biomedical applications. Our contributions can be summarized as follows:
  • By conditioning on 3D spatial gene expression as the input prompt, Tera-MIND enables the seamless generation of in-silico mouse brains at the scale of $0.77$ teravoxels. Moreover, our approach accurately preserves fine-grained morphological structures across cellular, tissue, and slice-wise scales, supporting detailed exploration and comparison of brain architecture.
  • Leveraging 3D gene-gene self-attention mechanisms, we quantify and visualize spatial molecular interactions of key pathways that contribute to fundamental brain functions, including those involved in glutamatergic (Slc17a6 and Slc17a7) and dopaminergic (Nr4a2 and Th) neuronal systems.
  • Similar to the translational aims of wet-lab mouse brain studies, we show that Tera-MIND’s simulation results are applicable to previously unseen human health and glioblastoma brain samples, underscoring its potential in advancing the understanding of complex brain functions and diseases for real-world biomedical applications.
RESULTS
Model overview of Tera-MIND
[…]
Tera-MIND accurately generated tera-scale mouse brain(s) by spatial gene expression.
Consistent with the primary analysis described in 2, which was conducted on a comprehensive brain atlas of a P56 female adult mouse (Sampleid: 638850), we present main simulation results of Tera-MIND based on the same brain atlas. The holistic 3D comparison between the ground truth (GT) and our generation result is illustrated in Fig. 1 (b) and Fig. 2 (a). Given 3D spatial gene expression data as an input, Tera-MIND achieves the faithful and comprehensive reconstruction of 3D brain morphology, as captured by a sequential stack of DAPI and PolyT-stained WSIs. Specifically, the generated mouse brain has the same spatial resolution of 0.77 teravoxels as GT. At the level of individual WSIs (Fig. 2 (b, c)), Tera-MIND exhibits high-quality generation across multiple scales, including cellular, regional, and slice-wise morphological organization. Within distinct brain regions, such as the olfactory areas, isocortex, and cerebellar cortex, Tera-MIND accurately reconstructs complex morphological structures while preserving native cellular distributions. These results highlight the robustness of Tera-MIND for reconstructing 3D brain morphology directly from spatial gene expression profiles.
Complementary to the side-by-side qualitative comparisons, we perform thorough quantitative analyses of the reconstruction quality by Tera-MIND. For a fair and systematic assessment of the biomedical-specific outputs, we report not only commonly used metrics such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), but also include patch-based morphometric analyses of nuclear size and cell number. These domain-specific metrics are critical for evaluating the biological fidelity of reconstructed brain morphology at the cellular level. Then, we perform quanlitative (Fig. 2 (d)) and quantitative (Fig. 2 (e)) evaluation on Tera-MIND in comparison to state-of-the-art (SOTA) methods CoCoGAN 20, InfinityGAN 21, IST-editing 25, SinFusion 23, Patch-DM 24, etc. Overall, Tera-MIND achieved superior performance in terms of better PSNR and SSIM scores. Furthermore, experimental results reported in Fig. 2 (e) demonstrated marginal morphometric discrepancies between GT cellular patches and those generated by Tera-MIND, which are consistent with the results of stratified patch examples shown in Fig. 2 (e). Using spatial gene expression as the input prompt, these analyses confirm that Tera-MIND reliably reconstructs brain morphology across multiple scales, from cellular and regional structures to whole-brain organization. This consistent performance across scales highlights the efficacy of our approach for accurately capturing the complex spatial patterns of 3D brain morphology.
Tera-MIND identified biologically relevant gene-gene interactions in native 3D space.
To explore the role of individual genes and their interactions in governing the reconstruction of the 3D mouse brain, we analyze the learned 3D gene-gene attention map and focus on two critical neuronal systems: glutamatergic (GLUT) and dopaminergic (DOPA) signaling pathways. These examples demonstrate that Tera-MIND is capable of capturing biologically relevant spatial molecular interactions that are essential for the structural and functional interplay of mammalian brains.
GLUT neuronal system
Glutamate is the primary excitatory neurotransmitter in the brain and plays a pivotal role in a wide range of brain functions 26. The vesicular glutamate transporters VGLUT1 and VGLUT2, encoded by Slc17a7 and Slc17a6 resp., regulate extracellular glutamate concentrations and synaptic signaling 26, 27. Moreover, the expression level of VGLUTs impacts the amount of glutamate loaded into vesicles and released, thereby affecting neurotransmission 28. Changes in VGLUT expression levels are associated with various neurological pathologies such as schizophrenia, neuropathic pain, and ischemia 29, 30. In adult mouse brains, spatial expression patterns of VGLUT1 and VGLUT2 are generally complementary but not exclusive  28. Specifically, Slc17a7 (VGLUT1) is broadly expressed in the cerebral cortex, cerebellar cortex, hippocampus, and thalamus, etc., while Slc17a6 (VGLUT2) can be found in subcortical regions such as the thalamus and spinal cord.
As shown in Fig. 3 (a, left) and (b), spatial expression patterns of these transporters are accurately mapped in 3D space and align with prior studies 1, 2, which justifies the registration process of brain atlas utilized in this study. Interestingly, the 3D gene-gene attention map (Fig. 3 (a, middle)) learned by Tera-MIND reveals widespread and heterogeneous attention signals (Fig. 3 (a, right)) for the spatial interaction between Slc17a6 and Slc17a7. Such heterogeneous attention levels can also be observed by the marginal distribution reported in Fig. 3 (c), where the linear regression analysis identifies a strong correlation between expression and attention levels. Since both transporters collectively contribute to the regulation of synaptic vesicle glutamate content in those overlapped regions, these heterogeneous spatial attention signals (Fig. 3 (c, d)) likely encapsulate their accumulative impact on regional brain morphology and neurotransmission.
DOPA neuronal system
In dopamine-mediated neuronal signaling, Nr4a2 (Nurr1) is an important transcription factor that is expressed in several regions of the central nervous system including the olfactory bulb 31, the substantia nigra (SN) and ventral tegmental area (VTA). Nr4a2 plays a pivotal role in dopaminergic neurons by regulating several key genes (e.g., Th) involved in dopamine synthesis, storage, and release. Th encodes tyrosine hydroxylase, the rate-limiting enzyme in dopamine synthesis 31. Both Nr4a2 and Th genes are present in areas such as SN, VTA, and olfactory bulb of the mouse brain 32, 33.
The co-localization patterns have been correctly captured in spatial visualizations presented in Fig. 3 (e, left) and (f), which are subsequently used to derive spatial attention levels using self-attention mechanism (Fig. 3 (e, middle)). Reduced Nr4a2 and Th expression are associated with Parkinson’s disease 34, 35, a common neurodegenerative disorder with loss of dopaminergic neurons in the SN, leading to dopamine deficiency and motor function impairments 34, 35. This has also been observed in mice after deletion of Nr4a2 in mature dopaminergic neurons 36. Though the linear regression analysis of Fig. 3 (g) illustrates overall lower attention levels than the ones reported for GLUT system, we observe highly aggregated Nr4a2-Th attention signals in the olfactory bulb, SN and VTA (Bottom of midbrain label in Fig. 3 (h)), where both genes play a central role in regulating dopaminergic functions. As shown in Fig. 3 (e, right) and (h), the 3D gene-gene attention map uncovers a spatially aligned interaction between Nr4a2 and Th that reinforces their roles in the molecular regulation of dopaminergic pathways and cellular anatomical structures.
Tera-MIND achieved reproducible and robust results on three tera-scale mouse brains.
In addition to the main mouse brain discussed above, two additional P56 adult mouse brains – one male and one female – were profiled in previous studies 1, 2 for supporting analyses, as illustrated in Fig. 4 (a) (middle and right). To rigorously evaluate the reproducibility and robustness of our findings, we extend the same simulation experiments conducted on the main mouse brain to both supporting mouse brains (supp (m) and supp (f)).
Morphology generation: As illustrated in Fig. 4 (a, b), the side-by-side visual comparisons between our generated and GT results demonstrate convincing reconstructions of brain morphology on all three tera-scale mouse brains. Consistent with the main generation results (Fig. 1 (b) and the left plots of Fig. 4 (a, b)), our supporting experiments highlight competitive and reliable 3D and slice-wise image generation quality, as elaborated by the holistic and multi-scale visualization shown in the middle and right plots of Fig. 4 (a, b). For quantitative evaluation, Tera-MIND is benchmarked against seven SOTA methods COCO-GAN 20, InfinityGAN 21, MS-PIE 22, SST-editing 37, IST-editing 25, SinFusion 23, and Patch-DM 24. Complementary to PSNR and SSIM scores, we incorporate the spatial Fréchet Inception Distance (sFID) 38, which prioritizes spatially relevant evaluations by rewarding image distributions that preserve coherent spatial structures. Across all three brain instances, Tera-MIND consistently outperforms competing methods, in terms of achieving better PSNR, SSIM, and sFID scores (Fig. 4 (c)). When evaluated on domain-specific metrics such as cell number and nuclear size (Fig. 4 (d, e)), Tera-MIND yields minimal distributional discrepancies relative to GT. This highlights its capacity to maintain high-fidelity not only in structural reconstruction but in the reproduction of biologically relevant features.
Pathway identification: In the context of the GLUT pathway, Fig. 5 (a, b) present a highly consistent and spatially comparable localization of Slc17a6 and Slc17a7 expression patterns across two independent supporting instances. These expression maps are accurately registered in 3D space and align well with the ones from our main results and previous studies 1, 2. Similarly as observed in the main results, Fig. 5 (c) and (d) illustrate that Tera-MIND is capable of learning strong and heterogeneous Slc17a6-Slc17a7 attention levels driven by the positively correlated gene expression levels. This suggests their synergistic and collective roles in regulating neurotransmission and cellular morphology across overlapped regions. Regarding the DOPA pathway, despite the absence of the olfactory bulb in the raw data of the supporting mouse brains, the gene expression patterns of Nr4a2 and Th exhibit a comparable emergence in the SN and VTA regions (See also Fig. 5 (e, f)). Consistent with main results, the linear regression analysis of Fig. 5 demonstrates the comparable correlation between gene expression and attention levels. Moreover, we witness reproducible and clear identification of Nr4a2-Th attention levels derived from Tera-MIND in the same regions for both supporting mouse brains (Fig. 5 (e, h)). These results confirm a consistent expression pattern of critical genes associated with dopaminergic function in these specific regions. Please see also the side-by-side video visualizations for more details.
Tera-MIND facilitated the translational application on unseen human brain samples.
Following the translational aims of mouse brain studies in neuroscience, we conduct generalization experiments in which Tera-MIND was trained exclusively on mouse brain data (Main setting) and tested on the human brain healthy (Fig. 6 (a)) and glioblastoma (Fig. 6 (b)) samples  [^1] https://www.10xgenomics.com/datasets/xenium-human-brain-preview-data-1-standard. As shown in the gene panel (c) of Fig. 6, we firstly determine the targeted genes that are present in both mouse brain atlases and human brain samples, which are then fed into the Tera-MIND model as the conditional input. After training with the patch-wise paired data from mouse brain atlases, we run the WSI generation using the gene expression arrays profiled on the healthy and glioblastoma human samples. Both qualitative (Fig. 6 (a, b)) and quantitative evaluations (Fig. 6 (d)) confirm that the performance of morphological simulation is generalizable from mouse to human samples, in terms of well-preserved tissue architecture and meaningful cellular correspondence to GT. Notably, Tera-MIND reported better PSNR and SSIM for the healthy brain sample than in the setting of Glioblastoma, suggesting a mild cross-species translational gap under non-physiological conditions.
When examining the gene expression patterns of GLUT pathway between healthy and glioblastoma cases, Fig. 6 (a) and (b) reveal distinct spatial distributions: Slc17a7 exhibits higher and more heterogeneous expression levels across the glioblastoma sample compared to the healthy sample, resulting in pronounced interaction signals between Slc17a6 and Slc17a7. Conversely, both gene expression and their attention levels for the healthy sample remain homogeneous. This phenomenon can be attributed to the observation that glioma cells often release glutamate, leading to excitotoxic neuronal death and creating space for tumor expansion 39.
DISCUSSION
In this study, we proposed a novel patch-based diffusion model Tera-MIND for high-fidelity simulation of mammalian brains. For three tera-scale mouse brain atlases, Tera-MIND achieved reproducible generation results and consistently identified spatial molecular interactions within GLUT and DOPA pathways. Beyond the comprehensive evaluation on murine brains, we established the in-silico translational applicability of Tera-MIND by generalizing to previously unseen human brain samples, including both healthy and glioblastoma tissues. Methodologically, Tera-MIND offers an efficient and scalable generative framework for modeling whole (animal) organisms. From the viewpoint of clinical application, Tera-MIND thereby provides a novel in-silico approach for comparative pathology 40, which enables a systematic and side-by-side analysis of murine and human tissues to assess the validity and translational relevance of animal models. Furthermore, Tera-MIND inherently facilitates in-silico interventions on targeted genes of interest that drive (disease) morphological transitions 25. This capability allows for the exploration of ‘what if’ causal questions for enhancing clinical treatment and diagnostics 8. Given the scope of this study, we plan to address the topic of simulated intervention in future work.
Limitations of the study
In addition, we acknowledge the limitations of Tera-MIND. The modeling paradigm is fundamentally driven by the molecular-to-morphology spatial associations, which, while powerful, represent a simplified hypothesis of the complex neurological processes. With the rapid emergence of diverse spatial omics technologies 41, future work will incorporate orthogonal modalities, such as spatial proteomics and epigenomics, to provide additional layers of validation and further distinguish true functional spatial relationships.
In conclusion, the development of virtual replicas of entire biological organisms using GenAI, referred to as generative digital twins (GDTs 8) and exemplified by Tera-MIND in this study, opens new avenues for high-throughput biomedical simulation with minimal ethical concerns. This methodology offers a cost-effective and scalable alternative for preclinical testing of therapeutic strategies. Lastly, the potentially wide applicability of GDTs on human samples aligns with the principles of animal welfare by supporting the replacement, reduction, and refinement (3R) of animal use in laboratory research, providing a promising algorithmic tool for ethically responsible biomedical innovation.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Jiqing Wu (Jiqing.Wu@unibas.ch).
Data and code availability
  • The processed mouse brain data has been deposited at brain image library under the database identifier ace-lot-now and are publicly available as of the date of publication. All other data reported in this paper will be shared by the lead contact upon request.
  • All original code has been deposited at Zenodo under the DOI 10.5281/zenodo.14826874 and is publicly available as of the date of publication.
  • Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
This research project is mainly supported through the core professorship funding of Prof. Dr. med. Viktor Koelzer, Medical Co-Director and Professor of Experimental Pathology, University Hospital of Basel.
AUTHOR CONTRIBUTIONS
Conceptualization, J.W. and V.H.K.; methodology, J.W.; investigation, J.W., I.B., V.H.K.; writing-–original draft, J.W., V.H.K.; writing-–review & editing, J.W., I.B., V.H.K., Y.L., E.K.; funding acquisition, V.H.K.; supervision, V.H.K..
DECLARATION OF INTERESTS
V.H.K. reports being an invited speaker for Sharing Progress in Cancer Care (SPCC) and Indica Labs; advisory board of Takeda; and sponsored research agreements with Roche and IAG, all unrelated to the current study. V.H.K. is a participant of patent applications in computational pathology outside ofthe submitted work.
DECLARATION OF GENERATIVE AI AND AI-ASSISTED TECHNOLOGIES
During the preparation of this work, the author(s) used ChatGPT in order to polish the content. After using this tool or service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
SUPPLEMENTAL INFORMATION INDEX
Figures S1-S5 and their legends in a PDF
MAIN FIGURE TITLES AND LEGENDS
Figure 1: Overview of the Tera-MIND architecture and workflows for virtual organ simulation (Main result). a. The 3D reference map of major brain regions (left) alongside their corresponding labels (right) for guiding the navigation of brain structures. b. The 3D visualization of the ground-truth (GT, left) and generated (right) mouse brain, as captured by a sequential stack of DAPI and PolyT-stained WSIs at the scale of $0.77\times 10^{12}$ voxels. The bottom part of generated and GT results overlay brain region maps, aiding in the navigation of the complex spatial brain architecture. c. The conceptual illustration of patch-based diffusion model training. Here, we train Tera-MIND with noisy DAPI and PolyT bioimage patches so that the model learns to output clean patches guided by paired spatial mRNA array. Apart from the standard reconstruction path (black arrows), we supply the boundary-aware path (gray arrows) to impose boundary consistency between generated neighboring patches. d. The conceptual illustration of patch-wise generating virtual mouse brain using Tera-MIND. At this stage, we only run through the boundary-aware path to seamlessly generate the tera-scale mouse brain.
Figure 2: The visual and quantitative comparison of mouse brain generation (Main result). a. The 3D map of major brain regions (left), GT (middle), and the generated mouse brain (right) of the atlas used in the primary analysis 1. b. The cell- and region-level visualization of GT mouse brain WSIs in selected brain slices. For (a) and (b), the 3D/2D reference maps of brain regions are overlaid to aid in the navigation of brain structures. c. The cell- and region-level visualization of generated mouse brain WSIs. d. The visual comparison of exemplar cellular regions ($1024\times 1024$) between Tera-MIND and state-of-the-art (SOTA) patch-based models. From left to right, the image stack displays individual DAPI, PolyT, merged channels, and nuclear masks obtained using Cellpose 42. e. Quantitative comparison of image generation quality between Tera-MIND and SOTA methods. This includes mean and standard deviation scores (error bar) for widely-used image quality metrics, such as PSNR and SSIM, alongside domain-specific scores like patch-based nuclear size and cell count. For a more in-depth analysis, we encourage readers to zoom in on the visual results for clearer comparisons.
Figure 3: The visualization of gene-gene interactions for GLUT and DOPA neuronal systems (Main result). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The linear regression analysis of gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left), the conceptual illustration of 3D gene-gene attention layer (middle), and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The trending line showing the relationship between gene expr. versus attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. Here, both SN and VTA are located towards the bottom of midbrain label. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
Figure 4: The visual and quantitative comparison for the main and supp mouse brain results. a. The holistic 3D comparison between the generated and GT mouse brain, with the brain region map provided on the left for structural guidance. Note that all three generated mouse brains have the same $0.77\times 10^{12}$ voxel scale. b. The cell- and region-level visualizations of WSIs for the side-by-side comparison between the generated and GT mouse brain. c. The quantitative comparison of generation quality between Tera-MIND and SOTA methods, as measured by PSNR, SSIM, and sFID. d. Compared to the GT results, the quantitative evaluation of cell numbers derived from Tera-MIND and SOTA methods. e. Compared to the GT results, the quantitative evaluation of nuclear size derived from Tera-MIND and SOTA methods.
Figure 5: The visualization of gene-gene interaction for GLUT and DOPA neuronal systems (Supp (m) and (f) results). a. The 3D visualization of Slc17a6 and Slc17a7 gene expression (left) and the 3D visualization of Slc17a6-Slc17a7 attention (right). b. The WSI visualization of Slc17a6 and Slc17a7 gene expression (expr.). c. The linear regression analysis of gene expr. and attention level for both Slc17a6 and Slc17a7. d. The WSI visualization of Slc17a6-Slc17a7 attention level. e. The 3D visualization of Nr4a2 and Th gene expression (left) and the 3D visualization of Nr4a2-Th attention (right). f. The WSI visualization of Nr4a2 and Th gene expression (expr.). g. The linear regression analysis of gene expr. and attention level for both Nr4a2 and Th. h. The WSI visualization of Nr4a2-Th attention level. Here, both SN and VTA are located towards the bottom of midbrain label. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for both GLUT and DOPA pathways.
Figure 6: The comprehensive generation, quantification, and gene expression statistics for the human brain experiments. a. The side-by-side comparison between the generated (Gen) and ground-truth (GT) WSIs of the human brain (Healthy), with a $27136\times 36864$ spatial resolution. The visualization of Slc17a6 and Slc17a7 gene expression and their spatial attention. b. The side-by-side comparison between the Gen and GT WSIs of the human brain (Glioblastoma), with a $23808\times 39680$ spatial resolution. The visualization of Slc17a6 and Slc17a7 gene expression and their spatial attention. c. The side-by-side comparison of gene expression level (mean) for the human (Healthy) and human (Glioblastoma). d. The quantitative comparison of generation quality for the human (Healthy) and human (Glioblastoma), as measured by PSNR, SSIM, nuclear size, and cell number. For a clearer visualization, gene expr. and attention levels are normalized to the range [0, 1] for the GLUT pathway.